OpenStudy (anonymous):

find the average rate of change of f from pi to 11pi/3. f(x) = cos(x/2)

OpenStudy (anonymous):

\[m_{avg}={{\cos({11\pi \over 6})-\cos({\pi \over 2})}\over {11\pi \over 6}-{\pi \over 2}} \] \[m={{\sqrt{3} \over 2}\over {4\pi \over 3}}\] \[m={3\sqrt{3} \over 8\pi}\]

OpenStudy (anonymous):

thanks